# Can a $\xi R\phi^{2}$ non-minimal coupling contribute to the non-relativistic gravitational potential?

### v1

@@ 3.2 The vacuum polarisation
$-i{\cal M}^{\rm 3.1}=\frac{3\mu^{4\epsilon}\kappa^{4}\Lambda\xi}{8}(-q^{2})\frac{\eta^{\mu^{\prime}\nu^{\prime}}}{q^{2}}P_{\mu^{\prime}\nu^{\prime}\alpha^{\prime}\beta^{\prime}}\eta^{\alpha^{\prime}\beta^{\prime}}\eta^{\lambda^{\prime}\rho^{\prime}}\eta^{\gamma^{\prime}\delta^{\prime}}P_{\lambda^{\prime}\rho^{\prime}\lambda\rho}P_{\gamma^{\prime}\delta^{\prime}\gamma\delta}$
$\times\left\{k_{2(\mu}k^{\prime}_{2\nu)}-\eta_{\mu\nu}(k_{2}\cdot k_{2}^{\prime}+m^{2}_{2})\right\}\int\frac{d^{d}l}{(2\pi)^{d}}\frac{\bar{\tau}^{\mu^{\prime\prime}\nu^{\prime\prime}\lambda\rho\gamma\delta}}{l^{2}(l-q)^{2}}\frac{1}{q^{2}}P_{\mu^{\prime\prime}\nu^{\prime\prime}}{}^{\mu\nu}$
$+3\mu^{4\epsilon}\kappa^{4}\Lambda\xi(-q^{2})\frac{\eta^{\mu^{\prime}\nu^{\prime}}}{q^{2}}P_{\mu^{\prime}\nu^{\prime}\alpha^{\prime}\beta^{\prime}}P^{\beta^{\prime}\gamma^{\prime}}{}_{\lambda\rho}P_{\gamma^{\prime}}{}^{\alpha^{\prime}}{}_{\gamma\delta}$
$\times\left\{k_{2(\mu}k^{\prime}_{2\nu)}-\eta_{\mu\nu}(k_{2}\cdot k_{2}^{\prime}+m^{2}_{2})\right\}\int\frac{d^{d}l}{(2\pi)^{d}}\frac{\bar{\tau}^{\mu^{\prime\prime}\nu^{\prime\prime}\lambda\rho\gamma\delta}}{l^{2}(l-q)^{2}}\frac{1}{q^{2}}P_{\mu^{\prime\prime}\nu^{\prime\prime}}{}^{\mu\nu}$
$-\frac{3\mu^{4\epsilon}\kappa^{4}\Lambda\xi}{4}(-q^{2})\frac{\eta^{\mu^{\prime}\nu^{\prime}}}{q^{2}}P_{\mu^{\prime}\nu^{\prime}\alpha^{\prime}\beta^{\prime}}\eta^{\alpha^{\prime}\beta^{\prime}}P_{\lambda^{\prime}\rho^{\prime}\lambda\rho}P^{\lambda^{\prime}\rho^{\prime}}{}_{\gamma\delta}$
$\times\left\{k_{2(\mu}k^{\prime}_{2\nu)}-\eta_{\mu\nu}(k_{2}\cdot k_{2}^{\prime}+m^{2}_{2})\right\}\int\frac{d^{d}l}{(2\pi)^{d}}\frac{\bar{\tau}^{\mu^{\prime\prime}\nu^{\prime\prime}\lambda\rho\gamma\delta}}{l^{2}(l-q)^{2}}\frac{1}{q^{2}}P_{\mu^{\prime\prime}\nu^{\prime\prime}}{}^{\mu\nu}$
$-\frac{3\mu^{4\epsilon}\kappa^{4}\Lambda\xi}{2}(-q^{2})\frac{\eta^{\mu^{\prime}\nu^{\prime}}}{q^{2}}P_{\mu^{\prime}\nu^{\prime}\alpha^{\prime}\beta^{\prime}}P^{\alpha^{\prime}\beta^{\prime}}{}_{\lambda\rho}\eta_{\lambda^{\prime}\rho^{\prime}}P^{\lambda^{\prime}\rho^{\prime}}{}_{\gamma\delta}$
$\times\left\{k_{2(\mu}k^{\prime}_{2\nu)}-\eta_{\mu\nu}(k_{2}\cdot k_{2}^{\prime}+m^{2}_{2})\right\}\int\frac{d^{d}l}{(2\pi)^{d}}\frac{\bar{\tau}^{\mu^{\prime\prime}\nu^{\prime\prime}\lambda\rho\gamma\delta}}{l^{2}(l-q)^{2}}\frac{1}{q^{2}}P_{\mu^{\prime\prime}\nu^{\prime\prime}}{}^{\mu\nu}=0,$ (15)
where $\bar{\tau}^{\mu\nu\lambda\rho\gamma\delta}$ is the three graviton vertex of 8, modulo the factor $i\kappa/2$. Likewise, if the $k_{2}-k_{2}^{\prime}$ is non-minimal and it is connected to the $\Lambda$-three graviton vertex (i.e., the third of 3), the contribution vanishes. Similarly, we find all the contributions for this class of diagrams vanish.
@@ 3.3 The fish diagram
$-i{\cal M}^{\rm 4.1}=-\Lambda\mu^{4\epsilon}\kappa^{4}\xi(-q)^{2}\eta^{\mu\nu}\frac{P_{\mu\nu\lambda^{\prime}\rho^{\prime}}}{q^{2}}\eta^{\lambda^{\prime}\rho^{\prime}}\int\frac{d^{d}l}{(2\pi)^{d}}\frac{\eta_{\alpha^{\prime}\beta^{\prime}}P^{\alpha^{\prime}\beta^{\prime}}{}_{\lambda\eta}\eta_{\rho^{\prime}\sigma^{\prime}}P^{\rho^{\prime}\sigma^{\prime}}{}_{\rho\sigma}}{l^{2}(l-q)^{2}}\times\bar{\tau}_{2,2}^{\lambda\eta\rho\sigma}$
$-8\mu^{4\epsilon}\Lambda\kappa^{4}\xi(-q)^{2}\eta^{\mu\nu}\frac{P_{\mu\nu\mu^{\prime}\nu^{\prime}}}{q^{2}}P^{\nu^{\prime}\lambda^{\prime}}{}_{\lambda\eta}P_{\lambda^{\prime}}{}^{\mu^{\prime}}_{\rho\sigma}\int\frac{d^{d}l}{(2\pi)^{d}}\frac{1}{l^{2}(l-q)^{2}}\times\bar{\tau}_{2,2}^{\lambda\eta\rho\sigma}$
$+2\mu^{4\epsilon}\Lambda\kappa^{4}\xi(-q)^{2}\eta^{\mu\nu}\frac{P_{\mu\nu\mu^{\prime}\nu^{\prime}}}{q^{2}}\eta^{\mu^{\prime}\nu^{\prime}}P^{\lambda^{\prime}\rho^{\prime}}{}_{\lambda\eta}P_{\lambda^{\prime}\rho^{\prime}\rho\sigma}\int\frac{d^{d}l}{(2\pi)^{d}}\frac{1}{l^{2}(l-q)^{2}}\times\bar{\tau}_{2,2}^{\lambda\eta\rho\sigma}$
$+4\mu^{4\epsilon}\Lambda\kappa^{4}\xi(-q)^{2}\eta^{\mu\nu}\frac{P_{\mu\nu\mu^{\prime}\nu^{\prime}}}{q^{2}}P^{\mu^{\prime}\nu^{\prime}}{}_{\lambda\eta}\eta^{\lambda^{\prime}\rho^{\prime}}P_{\lambda^{\prime}\rho^{\prime}\rho\sigma}\int\frac{d^{d}l}{(2\pi)^{d}}\frac{1}{l^{2}(l-q)^{2}}\times\bar{\tau}_{2,2}^{\lambda\eta\rho\sigma}$ (16)
@@
A little algebra shows that 16 is vanishing. Likewise if the ${\cal O}(\kappa)$ non-minimal vertex is $k_{2}-k^{\prime}_{2}$, we find vanishing contribution (the third of 4).
@@
$-i{\cal M}^{\rm 4.2}=\frac{\mu^{3\epsilon}\Lambda\kappa^{4}\xi}{4}k_{1(\mu}k^{\prime}_{1\nu)}P^{\mu\nu\mu^{\prime}\nu^{\prime}}\eta_{\mu^{\prime}\nu^{\prime}}\int\frac{d^{d}l}{(2\pi)^{d}}\frac{1}{l^{2}(l-q)^{2}q^{2}}$
@@
$\times\left\{-\frac{1}{2}\eta^{\lambda^{\prime}\rho^{\prime}}P_{\lambda^{\prime}\rho^{\prime}}{}^{\mu\alpha}\eta^{\alpha^{\prime}\beta^{\prime}}P_{\alpha^{\prime}\beta^{\prime}\mu\lambda}l^{\lambda}(l-q)_{\alpha}+\frac{3}{4}l\cdot(l-q)\eta^{\lambda^{\prime}\rho^{\prime}}P_{\lambda^{\prime}\rho^{\prime}}{}^{\mu\nu}\eta^{\alpha^{\prime}\beta^{\prime}}P_{\alpha^{\prime}\beta^{\prime}\mu\nu}\right\}$
$+2\mu^{3\epsilon}\Lambda\kappa^{4}\xi k_{1(\mu^{\prime}}k^{\prime}_{1\nu^{\prime})}P^{\mu^{\prime}\nu^{\prime}}{}_{\lambda^{\prime}\rho^{\prime}}\int\frac{d^{d}l}{(2\pi)^{d}}\frac{1}{l^{2}(l-q)^{2}q^{2}}$
$\times\left\{-\frac{1}{2}P^{\rho^{\prime}\gamma^{\prime}\mu\alpha}{P_{\gamma^{\prime}}{}^{\lambda^{\prime}}}_{\mu\lambda}l^{\lambda}(l-q)_{\alpha}+\frac{3}{4}l\cdot(l-q)P^{\gamma^{\prime}\rho^{\prime}\mu\nu}P^{\lambda^{\prime}}{}_{\gamma^{\prime}\mu\nu}\right\}$
$-\mu^{3\epsilon}\Lambda\kappa^{4}\xi k_{1(\mu^{\prime}}k^{\prime}_{1\nu^{\prime})}P^{\mu^{\prime}\nu^{\prime}\alpha^{\prime}\beta^{\prime}}\int\frac{d^{d}l}{(2\pi)^{d}}\frac{1}{l^{2}(l-q)^{2}q^{2}}$
$\times\left\{-\frac{1}{2}P_{\alpha^{\prime}\beta^{\prime}}{}^{\mu\alpha}\eta^{\lambda^{\prime}\rho^{\prime}}P_{\lambda^{\prime}\rho^{\prime}\mu\lambda}l^{\lambda}(l-q)_{\alpha}+\frac{3}{4}l\cdot(l-q)P_{\alpha^{\prime}\beta^{\prime}}{}^{\mu\nu}\eta^{\lambda^{\prime}\rho^{\prime}}P_{\lambda^{\prime}\rho^{\prime}\mu\nu}\right\}$
$-\mu^{3\epsilon}\frac{\Lambda\kappa^{4}\xi}{2}k_{1(\mu^{\prime}}k^{\prime}_{1\nu^{\prime})}P^{\mu^{\prime}\nu^{\prime}\alpha^{\prime}\beta^{\prime}}\eta_{\alpha^{\prime}\beta^{\prime}}\int\frac{d^{d}l}{(2\pi)^{d}}\frac{1}{l^{2}(l-q)^{2}q^{2}}$
$\times\left\{-\frac{1}{2}P_{\gamma\delta}{}^{\mu\alpha}P^{\gamma\delta}{}_{\mu\lambda}l^{\lambda}(l-q)_{\alpha}+\frac{3}{4}l\cdot(l-q)P_{\gamma\delta}{}^{\mu\nu}P^{\gamma\delta}{}_{\mu\nu}\right\}$
$=\frac{\mu^{3\epsilon}m_{1}^{2}\Lambda\kappa^{4}\xi}{4}\int\frac{d^{d}l}{(2\pi)^{d}}\frac{1}{l^{2}(l-q)^{2}}-\frac{\mu^{3\epsilon}\Lambda\kappa^{4}\xi}{q^{2}}\int\frac{d^{d}l}{(2\pi)^{d}}\frac{2k_{1\mu}k^{\prime}_{1\nu}l^{\mu}l^{\nu}+q^{2}l\cdot q/2}{l^{2}(l-q)^{2}}$ (18)
@@
Only the first integral contributes to the potential at leading order, which is $\sim\ln q^{2}$. Likewise by replacing $m_{1}$ by $m_{2}$, we obtain the result if $k_{2}-k^{\prime}_{2}$ is the non-minimal two scalar-two graviton vertex. Putting things together now including 14, we have the total Feynman amplitude for 2-2 scattering at ${\cal O}(\xi\Lambda\kappa^{4})$, in the non-relativistic limit,
$-i{\cal M}^{\rm tot.,non-analytic}|_{\rm NR}=-\frac{i\Lambda\kappa^{4}\xi m_{1}^{2}m_{2}^{2}}{2\pi^{2}\vec{q}^{2}}-\frac{i(m_{1}^{2}+m_{2}^{2})\Lambda\kappa^{4}\xi}{4\pi^{2}}\ln\frac{\vec{q}^{2}}{\mu^{2}}$ (19)

### v2

@@ 3.2 The vacuum polarisation
$\begin{split}-i{\cal M}^{\rm 3.1}=&\frac{3\mu^{4\epsilon}\kappa^{4}\Lambda\xi}{8}(-q^{2})\frac{\eta^{\mu^{\prime}\nu^{\prime}}}{q^{2}}P_{\mu^{\prime}\nu^{\prime}\alpha^{\prime}\beta^{\prime}}\eta^{\alpha^{\prime}\beta^{\prime}}\eta^{\lambda^{\prime}\rho^{\prime}}\eta^{\gamma^{\prime}\delta^{\prime}}P_{\lambda^{\prime}\rho^{\prime}\lambda\rho}P_{\gamma^{\prime}\delta^{\prime}\gamma\delta}\\ &\times\left\{k_{2(\mu}k^{\prime}_{2\nu)}-\eta_{\mu\nu}(k_{2}\cdot k_{2}^{\prime}+m^{2}_{2})\right\}\int\frac{d^{d}l}{(2\pi)^{d}}\frac{\bar{\tau}^{\mu^{\prime\prime}\nu^{\prime\prime}\lambda\rho\gamma\delta}}{l^{2}(l-q)^{2}}\frac{1}{q^{2}}P_{\mu^{\prime\prime}\nu^{\prime\prime}}{}^{\mu\nu}+3\mu^{4\epsilon}\kappa^{4}\Lambda\xi(-q^{2})\frac{\eta^{\mu^{\prime}\nu^{\prime}}}{q^{2}}P_{\mu^{\prime}\nu^{\prime}\alpha^{\prime}\beta^{\prime}}P^{\beta^{\prime}\gamma^{\prime}}{}_{\lambda\rho}P_{\gamma^{\prime}}{}^{\alpha^{\prime}}{}_{\gamma\delta}\\ &\times\left\{k_{2(\mu}k^{\prime}_{2\nu)}-\eta_{\mu\nu}(k_{2}\cdot k_{2}^{\prime}+m^{2}_{2})\right\}\int\frac{d^{d}l}{(2\pi)^{d}}\frac{\bar{\tau}^{\mu^{\prime\prime}\nu^{\prime\prime}\lambda\rho\gamma\delta}}{l^{2}(l-q)^{2}}\frac{1}{q^{2}}P_{\mu^{\prime\prime}\nu^{\prime\prime}}{}^{\mu\nu}-\frac{3\mu^{4\epsilon}\kappa^{4}\Lambda\xi}{4}(-q^{2})\frac{\eta^{\mu^{\prime}\nu^{\prime}}}{q^{2}}P_{\mu^{\prime}\nu^{\prime}\alpha^{\prime}\beta^{\prime}}\eta^{\alpha^{\prime}\beta^{\prime}}P_{\lambda^{\prime}\rho^{\prime}\lambda\rho}P^{\lambda^{\prime}\rho^{\prime}}{}_{\gamma\delta}\\ &\times\left\{k_{2(\mu}k^{\prime}_{2\nu)}-\eta_{\mu\nu}(k_{2}\cdot k_{2}^{\prime}+m^{2}_{2})\right\}\int\frac{d^{d}l}{(2\pi)^{d}}\frac{\bar{\tau}^{\mu^{\prime\prime}\nu^{\prime\prime}\lambda\rho\gamma\delta}}{l^{2}(l-q)^{2}}\frac{1}{q^{2}}P_{\mu^{\prime\prime}\nu^{\prime\prime}}{}^{\mu\nu}-\frac{3\mu^{4\epsilon}\kappa^{4}\Lambda\xi}{2}(-q^{2})\frac{\eta^{\mu^{\prime}\nu^{\prime}}}{q^{2}}P_{\mu^{\prime}\nu^{\prime}\alpha^{\prime}\beta^{\prime}}P^{\alpha^{\prime}\beta^{\prime}}{}_{\lambda\rho}\eta_{\lambda^{\prime}\rho^{\prime}}P^{\lambda^{\prime}\rho^{\prime}}{}_{\gamma\delta}\\ &\times\left\{k_{2(\mu}k^{\prime}_{2\nu)}-\eta_{\mu\nu}(k_{2}\cdot k_{2}^{\prime}+m^{2}_{2})\right\}\int\frac{d^{d}l}{(2\pi)^{d}}\frac{\bar{\tau}^{\mu^{\prime\prime}\nu^{\prime\prime}\lambda\rho\gamma\delta}}{l^{2}(l-q)^{2}}\frac{1}{q^{2}}P_{\mu^{\prime\prime}\nu^{\prime\prime}}{}^{\mu\nu}=0\end{split}$ (17)
where $\bar{\tau}^{\mu\nu\lambda\rho\gamma\delta}$ is the three graviton vertex of 8, modulo the factor $i\kappa/2$. Likewise, if the $k_{2}-k_{2}^{\prime}$ is non-minimal and it is connected to the $\Lambda$-three graviton vertex (i.e., the third of 3), the contribution vanishes. Similarly, we find that all the contributions for this class of diagrams vanish.
@@ 3.3 The fish diagram
$\begin{split}-i{\cal M}^{\rm 4.1}=&-\Lambda\mu^{4\epsilon}\kappa^{4}\xi(-q)^{2}\eta^{\mu\nu}\frac{P_{\mu\nu\lambda^{\prime}\rho^{\prime}}}{q^{2}}\eta^{\lambda^{\prime}\rho^{\prime}}\int\frac{d^{d}l}{(2\pi)^{d}}\frac{\eta_{\alpha^{\prime}\beta^{\prime}}P^{\alpha^{\prime}\beta^{\prime}}{}_{\lambda\eta}\eta_{\rho^{\prime}\sigma^{\prime}}P^{\rho^{\prime}\sigma^{\prime}}{}_{\rho\sigma}}{l^{2}(l-q)^{2}}\times\bar{\tau}_{2,2}^{\lambda\eta\rho\sigma}\\ &-8\mu^{4\epsilon}\Lambda\kappa^{4}\xi(-q)^{2}\eta^{\mu\nu}\frac{P_{\mu\nu\mu^{\prime}\nu^{\prime}}}{q^{2}}P^{\nu^{\prime}\lambda^{\prime}}{}_{\lambda\eta}P_{\lambda^{\prime}}{}^{\mu^{\prime}}_{\rho\sigma}\int\frac{d^{d}l}{(2\pi)^{d}}\frac{1}{l^{2}(l-q)^{2}}\times\bar{\tau}_{2,2}^{\lambda\eta\rho\sigma}\end{split}$ (18)
$\begin{split}&+2\mu^{4\epsilon}\Lambda\kappa^{4}\xi(-q)^{2}\eta^{\mu\nu}\frac{P_{\mu\nu\mu^{\prime}\nu^{\prime}}}{q^{2}}\eta^{\mu^{\prime}\nu^{\prime}}P^{\lambda^{\prime}\rho^{\prime}}{}_{\lambda\eta}P_{\lambda^{\prime}\rho^{\prime}\rho\sigma}\int\frac{d^{d}l}{(2\pi)^{d}}\frac{1}{l^{2}(l-q)^{2}}\times\bar{\tau}_{2,2}^{\lambda\eta\rho\sigma}\\ &+4\mu^{4\epsilon}\Lambda\kappa^{4}\xi(-q)^{2}\eta^{\mu\nu}\frac{P_{\mu\nu\mu^{\prime}\nu^{\prime}}}{q^{2}}P^{\mu^{\prime}\nu^{\prime}}{}_{\lambda\eta}\eta^{\lambda^{\prime}\rho^{\prime}}P_{\lambda^{\prime}\rho^{\prime}\rho\sigma}\int\frac{d^{d}l}{(2\pi)^{d}}\frac{1}{l^{2}(l-q)^{2}}\times\bar{\tau}_{2,2}^{\lambda\eta\rho\sigma}\end{split}$
@@
A little algebra shows that 18 is vanishing. Likewise, if the ${\cal O}(\kappa)$ non-minimal vertex is $k_{2}-k^{\prime}_{2}$, we find a vanishing contribution (the third of 4).
@@
$\begin{split}-i{\cal M}^{\rm 4.2}=&\frac{\mu^{3\epsilon}\Lambda\kappa^{4}\xi}{4}k_{1(\mu}k^{\prime}_{1\nu)}P^{\mu\nu\mu^{\prime}\nu^{\prime}}\eta_{\mu^{\prime}\nu^{\prime}}\int\frac{d^{d}l}{(2\pi)^{d}}\frac{1}{l^{2}(l-q)^{2}q^{2}}\\ &\times\left\{-\frac{1}{2}\eta^{\lambda^{\prime}\rho^{\prime}}P_{\lambda^{\prime}\rho^{\prime}}{}^{\mu\alpha}\eta^{\alpha^{\prime}\beta^{\prime}}P_{\alpha^{\prime}\beta^{\prime}\mu\lambda}l^{\lambda}(l-q)_{\alpha}+\frac{3}{4}l\cdot(l-q)\eta^{\lambda^{\prime}\rho^{\prime}}P_{\lambda^{\prime}\rho^{\prime}}{}^{\mu\nu}\eta^{\alpha^{\prime}\beta^{\prime}}P_{\alpha^{\prime}\beta^{\prime}\mu\nu}\right\}\\ &+2\mu^{3\epsilon}\Lambda\kappa^{4}\xi k_{1(\mu^{\prime}}k^{\prime}_{1\nu^{\prime})}P^{\mu^{\prime}\nu^{\prime}}{}_{\lambda^{\prime}\rho^{\prime}}\int\frac{d^{d}l}{(2\pi)^{d}}\frac{1}{l^{2}(l-q)^{2}q^{2}}\left\{-\frac{1}{2}P^{\rho^{\prime}\gamma^{\prime}\mu\alpha}{P_{\gamma^{\prime}}{}^{\lambda^{\prime}}}_{\mu\lambda}l^{\lambda}(l-q)_{\alpha}+\frac{3}{4}l\cdot(l-q)P^{\gamma^{\prime}\rho^{\prime}\mu\nu}P^{\lambda^{\prime}}{}_{\gamma^{\prime}\mu\nu}\right\}\\ &-\mu^{3\epsilon}\Lambda\kappa^{4}\xi k_{1(\mu^{\prime}}k^{\prime}_{1\nu^{\prime})}P^{\mu^{\prime}\nu^{\prime}\alpha^{\prime}\beta^{\prime}}\int\frac{d^{d}l}{(2\pi)^{d}}\frac{1}{l^{2}(l-q)^{2}q^{2}}\left\{-\frac{1}{2}P_{\alpha^{\prime}\beta^{\prime}}{}^{\mu\alpha}\eta^{\lambda^{\prime}\rho^{\prime}}P_{\lambda^{\prime}\rho^{\prime}\mu\lambda}l^{\lambda}(l-q)_{\alpha}+\frac{3}{4}l\cdot(l-q)P_{\alpha^{\prime}\beta^{\prime}}{}^{\mu\nu}\eta^{\lambda^{\prime}\rho^{\prime}}P_{\lambda^{\prime}\rho^{\prime}\mu\nu}\right\}\\ &-\mu^{3\epsilon}\frac{\Lambda\kappa^{4}\xi}{2}k_{1(\mu^{\prime}}k^{\prime}_{1\nu^{\prime})}P^{\mu^{\prime}\nu^{\prime}\alpha^{\prime}\beta^{\prime}}\eta_{\alpha^{\prime}\beta^{\prime}}\int\frac{d^{d}l}{(2\pi)^{d}}\frac{1}{l^{2}(l-q)^{2}q^{2}}\left\{-\frac{1}{2}P_{\gamma\delta}{}^{\mu\alpha}P^{\gamma\delta}{}_{\mu\lambda}l^{\lambda}(l-q)_{\alpha}+\frac{3}{4}l\cdot(l-q)P_{\gamma\delta}{}^{\mu\nu}P^{\gamma\delta}{}_{\mu\nu}\right\}\\ =&\frac{\mu^{3\epsilon}m_{1}^{2}\Lambda\kappa^{4}\xi}{4}\int\frac{d^{d}l}{(2\pi)^{d}}\frac{1}{l^{2}(l-q)^{2}}-\frac{\mu^{3\epsilon}\Lambda\kappa^{4}\xi}{q^{2}}\int\frac{d^{d}l}{(2\pi)^{d}}\frac{2k_{1\mu}k^{\prime}_{1\nu}l^{\mu}l^{\nu}+q^{2}l\cdot q/2}{l^{2}(l-q)^{2}}\end{split}$ (20)
@@
Only the first integral contributes to the potential at leading order, which is $\sim\ln q^{2}$. Likewise, by replacing $m_{1}$ by $m_{2}$, we obtain the result if $k_{2}-k^{\prime}_{2}$ is the non-minimal two scalar-two graviton vertex. Putting things together now, including 16, we have the total Feynman amplitude for 2-2 scattering at ${\cal O}(\xi\Lambda\kappa^{4})$, in the non-relativistic limit,
$-i{\cal M}^{\rm tot.,non-analytic}|_{\rm NR}=-\frac{i\Lambda\kappa^{4}\xi m_{1}^{2}m_{2}^{2}}{2\pi^{2}\vec{q}^{2}}-\frac{i(m_{1}^{2}+m_{2}^{2})\Lambda\kappa^{4}\xi}{4\pi^{2}}\ln\frac{\vec{q}^{2}}{\mu^{2}}+\ {\cal O}\left(\vec{q}^{2}\ln\frac{\vec{q}^{2}}{\mu^{2}}\right).$ (21)
Note that since $|\vec{q}|$ is non-relativistic, we must have $|\vec{q}|/m\ll 1$. We expect the contribution from the first term in the potential to dominate the other.
Before we compute the potential corresponding to 21, a couple of comments on vanishing of some Feynman amplitudes are in order. Specifically, we have seen that two of the seagull diagrams falling into the topology of the first of 2 vanish. Likewise the amplitudes for all the vacuum polarisation diagrams, 3, vanish, whereas for 4, only two of the diagrams give non-vanishing contributions. Such vanishing amplitudes might correspond to some symmetry associated with the non-minimal and/or the $\Lambda$-three graviton vertices. For all these cases, note that the diagrams are nicely symmetric with respect to the position of the non-minimal vertex. We also note that two of the diagrams 4, even though they look symmetric with respect to the location of the non-minimal vertex, the said vertex is two scalar-two graviton and not two scalar-one graviton, and the diagrams yield non-vanishing contributions, 20. Thus the aforesaid symmetry, if there is any, seems to be only relevant to the non-minimal matter-graviton 3-point vertex, along with the $\Lambda$-three graviton vertex. In particular, from the structure of 17 it is not difficult to see that if we compute the one loop self energy of the graviton at $\Lambda\kappa^{2}$} due to one $\Lambda$-three graviton vertex and one usual three graviton vertex, the self energy would vanish. This corresponds to the cancellation between various terms arising due to the $\Lambda$-vertex. However, any specific nature or form of this speculated symmetry remains elusive to us so far. In this regard, we also refer our reader to e.g. [51] for demonstration of vanishing of one loop one and two point functions in a theory of inflation containing higher derivative terms for the inflaton. We also refer to [52] for vanishing of some scattering amplitudes off the Kerr black hole.